For a certain chemical reaction, the standard Gibbs free energy of reaction is −141. kJ. Calculate the temperature at which the equilibrium constant K=3.6 x 10^25.
Answer #1

The relation between Gibb's free energy and the equilibrium constant is given by the following equation :

G = - RT lnKeq

Therefore, - 141*1000 J = - 8.314 JK-1mol-1*T*(ln 3.6*1025)

Therefore, T= 141000/8.314 K-1mol-1(ln 3.6*1025)

Therefore, T= 16959 K/58.85= 288 K

Therefore, the temperature at the given conditions is 288 K.

Thank You!